# PrPSeg: Universal Proposition Learning for Panoramic Renal Pathology Segmentation

### 29 Feb 2024 v1

@@ 1 Introduction
To address these challenges, we introduce a token-based dynamic head network designed to achieve panoramic renal pathology segmentation (PrPSeg) by modeling the spatial relationships among all objects. This approach allows for the reuse of the same architectural framework, even when the dataset size expands. A universal proposition matrix is established to translate anatomical relationships into computational modeling concepts. A novel anatomy loss is also proposed to integrate spatial relationships into the model training as a form of semi-supervised learning. To our knowledge, this is the first deep learning algorithm to accomplish panoramic segmentation in renal pathology, demonstrating superior performance in all-encompassing segmentation.
@@ 2.1 Renal Pathology Segmentation
Building upon these insights, our work introduces a token-based approach, leveraging class-specific and scale-specific tokens. This method is designed to capture heterogeneous features and employs semi-supervised learning to understand spatial relationships across multiple scales, achieving panoramic segmentation in renal pathology.
@@ 2.2 Dynamic Single Network
While multi-head single network designs have been proposed for multi-class renal pathology segmentation [19, 13, 6, 8], they often require dense multi-class annotations. Given the labor-intensive nature of such annotations, pathology data is frequently partially labeled. Additionally, forming spatial correlations, such as subset/superset relationships, remains a challenge in multi-class segmentation frameworks.
@@ 3.2 Universal proposition matrix
To translate these anatomy concepts into an engineering framework, we adopt Aristotle’s logic theory to develop a anatomical map for renal pathology. Aristotle’s theory examines the relationships between objects using four fundamental categorical propositions. Upon closer inspection, complex propositions reveal themselves as collections of simpler claims derived from these initial propositions. Specifically, we utilize two terms from Aristotle’s theory: (1) Universal Affirmation: “All S are P,” and (2) Universal Negation: “No S are P,”, to universally assert properties for all group members, indicating strong constraints and relationships. For example, in the context of two kidney structures, $A$ and $B$, if $B$ is within $A$, $B$ is a subset of $A$, following the rule of Universal Affirmation. Conversely, if $A$ and $B$ have no inclusion relationship, they are mutually exclusive, aligning with Universal Negation. These propositions are employed to construct a anatomical relationship map representing the classification and spatial relationships among renal pathology objects, as illustrated in Figure 2.
@@ 3.4 Anatomy Loss Function
$m=M(i,j)$ (6)
$L_{\text{hats}}(i,j)=$ $\frac{(m+1)(m-2)}{-2}\cdot\text{DCE}(1-Y_{i},Y^{\prime}_{j})$ (7)
$-\frac{(m-1)(m-2)}{6}\cdot\text{DCE}(Y_{i},Y_{i}\cup Y^{\prime}_{j})$
$+\frac{(m-1)(m+1)}{3}\cdot\text{DCE}(Y_{i},Y^{\prime}_{j})$
where $M$ is the anatomy matrix and DCE denotes the Dice Loss. The total loss function is an aggregate of supervised and semi-supervised losses, weighted by $\lambda_{\text{hats}}$.
@@
$+\lambda_{\text{hats}}\sum_{j=1}^{n}L_{\text{hats}}(i,j)\quad(j\neq i)$
@@ 4.2 Experiment Details
The training process of our model was divided into two distinct phases. In the initial phase, spanning the first 50 epochs, we employed a supervised learning strategy focused on minimizing binary dice loss and cross-entropy loss. Subsequently, for the remaining epochs, both supervised and semi-supervised learning strategies were utilized, incorporating anatomy loss to explore the spatial correlation among multiple objects. In our experiments, $\lambda_{\text{hats}}$ is 0.1.

### 21 Mar 2024 v2

@@ 1 Introduction
To address these challenges, we introduce a token-based dynamic head network designed to achieve panoramic renal pathology segmentation (PrPSeg) by modeling the spatial relationships among all objects. This approach allows for the reuse of the same architectural framework, even when the dataset size expands. A universal proposition matrix is established to translate anatomical relationships into computational modeling concepts. A anatomy loss is also proposed to integrate spatial relationships into the model training as a form of semi-supervised learning. To our knowledge, this is the first deep learning algorithm to accomplish panoramic segmentation in renal pathology, demonstrating superior performance in all-encompassing segmentation.
@@ 2.1 Renal Pathology Segmentation
Recent methods utilize hierarchical information for semantic segmentation [36, 33] or classification and prediction [8]. However, these methods primarily focus on class-based relationships between objects. While all objects have a uniform resolution in natural images, this approach neglects the emphasis on pixel-wise anatomical and spatial relationships at multiple resolutions in kidney datasets.
Building upon these insights, our work introduces a token-based approach, leveraging class-specific and scale-specific tokens. This method is designed to capture heterogeneous features and employs semi-supervised learning to understand pixel-wise spatial relationships across multiple scales, achieving panoramic segmentation in renal pathology.
@@ 2.2 Dynamic Single Network
While multi-head single network designs have been proposed for multi-class renal pathology segmentation [20, 14, 6, 9], they often require dense multi-class annotations. Given the labor-intensive nature of such annotations, pathology data is frequently partially labeled. Additionally, forming spatial correlations, such as subset/superset relationships, remains a challenge in multi-class uplsegmentation frameworks.
@@ 3.2 Universal proposition matrix
To translate these anatomy concepts into an engineering framework, we adopt Aristotle’s logic theory to develop an anatomical map for renal pathology. Aristotle’s theory examines the relationships between objects using four fundamental categorical propositions. Upon closer inspection, complex propositions reveal themselves as collections of simpler claims derived from these initial propositions. Specifically, we utilize two terms from Aristotle’s theory: (1) Universal Affirmation: “All S are P," and (2) Universal Negation: “No S are P,", to universally assert properties for all group members, indicating strong constraints and relationships. For example, in the context of two kidney structures, $A$ and $B$, if $B$ is within $A$, $B$ is a subset of $A$, following the rule of Universal Affirmation. Conversely, if $A$ and $B$ have no inclusion relationship, they are mutually exclusive, aligning with Universal Negation. These propositions are employed to construct an anatomical relationship map representing the classification and spatial relationships among renal pathology objects, as illustrated in Figure 2.
@@ 3.4 Anatomy Loss Function
$L_{\text{upl}}(i,j)=\begin{cases}\text{DCE}(1-Y_{i},Y^{\prime}_{j}),\quad\text{if}\quad M_{t}(i,j)=1\\ -\text{DCE}(Y_{i},Y_{i}\cup Y^{\prime}_{j}),\quad\text{if}\quad M_{t}(i,j)=-1\\ \text{DCE}(Y_{i},Y^{\prime}_{j}),\quad\text{if}\quad M_{t}(i,j)=2\\ 0,\quad\text{if}\quad M_{t}(i,j)=0\\ \end{cases}$ (6)
where $M_{t}$ is the anatomy matrix and DCE denotes the Dice Loss. The total loss function is an aggregate of supervised and semi-supervised losses, weighted by $\lambda_{\text{upl}}$.
@@
$+\lambda_{\text{upl}}\sum_{j=1}^{n}L_{\text{upl}}(i,j)\quad(j\neq i)$
@@ 4.2 Experiment Details
The training process of our model was divided into two distinct phases. In the initial phase, spanning the first 50 epochs, we employed a supervised learning strategy focused on minimizing binary dice loss and cross-entropy loss. Subsequently, for the remaining epochs, both supervised and semi-supervised learning strategies were utilized, incorporating anatomy loss to explore the spatial correlation among multiple objects. In our experiments, $\lambda_{\text{upl}}$ is 0.1.
@@ 5.2 Ablation Study
We also provide an ablation study for two data extension scenarios in the Appendix (A.1). The proposed PrPSeg method is flexible to extend to new classes by merely updating tokens and the adaptable proposition matrix, without changing the backbone network, while demonstrating superior performance compared to baseline methods across all seven new classes.
@@ 7 Acknowledgement
Appendix A Appendix
A.1 Evaluation with Classes Extension
We provide an ablation study for two data extension scenarios: (1) adding 3 new sub-types; (2) introducing 4 new objects to our dataset. The proposed PrPSeg method is flexible to extend new classes by merely updating tokens and the adaptable proposition matrix, without changing the backbone network (Section A.2). All models are trained for 30 epochs on the 15-class dataset using the same codebase and experimental settings as those described in the main manuscript. In Table S4, PrPSeg demonstrated superior performance compared to baseline methods across all seven new classes, maintaining the trend observed in Table 3 of the manuscript.
Table S4: Ablation study with 7 extended classes. Dice similarity coefficient scores (%) are reported.
Method TDH UPL Regions Functional units Cells Mean Inn. Cor. Mid. Cor. Out. Cor. Art. PTC MV Smooth. Swin-Unetr 34.54 31.26 34.20 53.92 57.46 55.03 60.18 49.66 Swin-Unetr 45.25 41.89 70.22 52.27 60.95 52.17 62.45 56.67 Omni-Seg 39.84 43.98 70.96 47.33 63.23 48.67 56.91 53.86 Omni-Seg 51.16 46.46 70.53 57.89 62.63 61.93 64.41 60.43 PrPSeg (Ours) 52.69 49.86 71.13 59.51 64.74 63.09 64.91 61.74
*TDH is Token-based Dynamic Head *UPL is Universal Proposition Learning
A.2 Novelty Clarification
The contributions of this paper are threefold: (1) A comprehensive universal proposition matrix is proposed to provide a simple and adaptable method to model the predominantly overlooked intricate spatial interrelations and class relationships among objects from clinical knowledge. This proposition matrix allows us to flexibly add unseen new classes via minimal changes (only modify tokens and this matrix); (2) The development of a token-based dynamic head in a single network architecture, improving partial label image segmentation. The backbone of the proposed PrPSeg network remains unchanged when new class tokens are introduced for new datasets, enabling the reuse of model weights on incomplete datasets. (3) The formulation of an anatomical loss function that quantifies the inter-object relationships across the kidney.
Figure S6: The innovation of the pipeline when data is extended – The proposed PrPSeg method is flexible to extend new classes by merely updating tokens and the adaptable proposition matrix, without changing the backbone network.
A.3 Comparison to Related Work and Contributions Beyond Medical Imaging
Several recent methods that utilize hierarchical information for semantic segmentation [36, 33] or classificationand prediction [8]. Our method’s innovations, beyond previous work, include: (1) Emphasizing pixel-wise anatomical and spatial relationships between objects, rather than solely taxonomy-based relationships (e.g., a glomerulus is located inside the cortex, not merely as a subset or sharing the cortex’s morphology); (2) Introducing a hierarchical relationship with class tokens and scale tokens across multiple resolutions (regions at 5$\times$, cells at 20$\times$) to provide greater flexibility between classes and scales, as opposed to a uniform resolution in natural images; (3) Enhancing the extensibility and reusability in model design for data expansion. The proposed method aims to provide a pipeline that is scale-aware, adaptive, and anatomically aware, transitioning from the clinical domain to potential applications in incremental learning and multi-view, multi-scale learning beyond the medical field.